OpenStudy (anonymous):

how do u find the exact value of the expression: tan inverse[tan(-pi/4)]?

OpenStudy (anonymous):

arctan(tan(x))=x You're undoing what tan does to x. So here it would be -Pi/4

OpenStudy (anonymous):

not... tan−1(tanx)=x only for x (−π/2,π/2)

OpenStudy (anonymous):

I'm just asking

OpenStudy (anonymous):

i thought its -pi/2 please explain

OpenStudy (anonymous):

tan(x)=y/x tan(-Pi/4)=-1/1=-1 arctan(x) means tangent of what angle gives x? arctan(-1)=-Pi/4, or also 3Pi/4 is a solution

OpenStudy (anonymous):

shouldve said tan(theta)=y/x